1. -20 points SerCP9 24.P.001.W My Notes Ask Your Teacher A laser beam is incident on two slits with a separation of 0.230 mm, and a screen is placed 4.90 m from the slits. If the bright interference fringes on the screen are separated by 1.55 cm, what is the wavelength of the laser light? nm Need Help? Watch It Talk to a Tutor
Answer #1

Δx = λL / d → λ = d Δx / L

A laser beam is incident on two slits with a separation of 0.230mm (d)
A screen is placed 4.90m from the slits (L)
The bright interference fringes on the screen are separated by 1.55cm (Δx)

You have → 0.0155m = λ(4.90m) / (0.0002m)

What is the wavelength of the laser light?

λ = d Δx / L = (0.0002m)(0.0155m) / (4.90m) = 632nm